# Enhancing Reasoning for Diffusion LLMs via Distribution Matching Policy Optimization

### v1

@@ 3.3 Effective Training with Negative Gradient Insertion
We note that this issue does not arise when the batch size is sufficiently large for the following reason. When having a large batch of diverse responses that make up a good coverage of the sample space, despite having all positive weights, since the model cannot increase likelihood on all responses (as it is a probability distribution that sums up to $1$), the “bad” responses will be automatically and implicitly penalized due to not having larger weights than the “good” responses.
@@
When the batch size is small, the scenario is different as is illustrated in Fig.˜3. In such a case, the model will tend to promote both “good” and “bad” responses due to the positive weights, and potentially penalize the likelihood of other unseen responses to maintain a valid probability distribution. This could be detrimental to achieving distribution matching, as these unseen responses may have high reward values and correspond to an undiscovered distribution mode.
@@
To address this issue, we inject negative gradient (Ren & Sutherland, 2025; Deng et al., 2025) by designing a weight baseline and subtract it from the obtained weights to facilitate an effective reinforcement on the good samples, i.e.,
$w_{(}{\bm{o}}|{\bm{q}};{\bm{\sigma}})=w({\bm{o}}|{\bm{q}};{\bm{\sigma}})-w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ (12)
@@
Group weight baseline. When the dLLM policy is close to optimal, the original log weight $\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ should behave approximately like constants for a group of different responses $\{{\bm{o}}^{(n)}\}_{1\leq n\leq N}$, leading to nearly uniform weight value for $\{w({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})\}_{1\leq n\leq N}$ after group softmax. We can thus choose the baseline as $1$ to encourage convergence to this optimal situation:
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=1,\;\forall n.$ (13)
Individual weight baseline. We can also consider the individual weight value of each response. For samples with smaller log weights, a stronger penalization is more desirable. A natural, adaptive way of designing penalization strength is to use softmax over the log weights with negative reward: let $\ell_{-}({\bm{o}}|{\bm{q}};{\bm{\sigma}}):=-\frac{r({\bm{q}},{\bm{o}})}{\alpha}+\log\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$, and define
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=\frac{N\exp(\ell_{-}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)}))}{\sum_{k}\exp(\ell_{-}({\bm{o}}^{(k)}|{\bm{q}};{\bm{\sigma}}^{(k)}))},\;\forall n.$ (14)
Note that this $w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ now corresponds to a bad target distribution given by $p_{*-}({\bm{o}}|{\bm{q}})\propto_{\bm{o}}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{-r({\bm{q}},{\bm{o}})/\alpha}$ which is tilted by the negative reward. The minus sign in the loss before $w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ means we want to drive the dLLM policy away from this bad distribution.
@@
d1-LLaDA $\mathbf{75.28}$ 81.40 84.38 30.00 36.60 40.80 34.38 $26.56$ 30.47 21.97 $11.04$ $8.69$
DMPO-LLaDA (Ours) 74.83 $\mathbf{82.41}$ $\mathbf{85.22}$ 30.00 $\mathbf{38.20}$ $\mathbf{42.80}$ $\mathbf{64.84}$ $\mathbf{75.78}$ $\mathbf{79.69}$ $\mathbf{32.76}$ $\mathbf{24.56}$ $\mathbf{19.97}$
Model weight baseline. Finally, we can determine whether to promote or penalize specific responses by comparing $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ with the importance weight under the current model policy $p_{\theta}({\bm{o}}|{\bm{q}})$. This pushes the model further towards the optimal one $p_{*}({\bm{o}}|{\bm{q}})$. Note that this does not incur additional computation overhead as we can estimate $\log p_{{\bar{\theta}}}({\bm{o}}|{\bm{q}})$ using negative ELBO ˜2, which is already computed in the WDCE loss. Let $\ell_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}}):=\log\frac{p_{{\bar{\theta}}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$, and define
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=\frac{N\exp(\ell_{\theta}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)}))}{\sum_{k}\exp(\ell_{\theta}({\bm{o}}^{(k)}|{\bm{q}};{\bm{\sigma}}^{(k)}))},~\forall n.$ (15)
We remark that the group weight and model weight baselines ˜13 and 15 can also be interpreted as an approximate variance reduction. See Sec.˜B.2 for discussion.
@@ 3.4 Weighted Direct Discriminative Optimization
To explore the full potential of the distribution matching framework in ˜6, we also investigate other choices for the potential $\mathcal{F}$ other than the cross-entropy. One particularly interesting objective is the following direct discriminative optimization (DDO) loss,
$\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=-\operatorname{\mathbb{E}}_{p_{*}({\bm{o}}|{\bm{q}})}\log\sigma\left(\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\right)-\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}})}\log\sigma\left(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\right),$ (16)
where $\sigma(t)=1/(1+\mathrm{e}^{-t})$. The global optimum of ˜16 is also $p_{*}(\cdot|{\bm{q}})$, thus being a valid functional for distribution matching learning. For a more detailed justification, see Sec.˜B.3.
This is inspired by Zheng et al. (2025e), which proposes a GAN-like (Goodfellow et al., 2014) training loss for the SFT of vision models. One interesting trait of this objective is its natural incorporation of negative gradients for bad samples due to the GAN nature, as is shown in the analysis therein:
$\nabla_{\theta}\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=\sum_{\bm{o}}\sigma\left(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\right)(p_{\theta}({\bm{o}}|{\bm{q}})-p_{*}({\bm{o}}|{\bm{q}}))\nabla_{\theta}\log p_{\theta}({\bm{o}}|{\bm{q}}).$
From the expression, as the first term is always non-negative, and the middle term $p_{\theta}({\bm{o}}|{\bm{q}})-p_{*}({\bm{o}}|{\bm{q}})$ applies a penalty for bad response $\bm{o}$}, thus providing a gradient direction for increasing $p_{\theta}({\bm{o}}|{\bm{q}})$. Leveraging this property, we adapt it for RL finetuning of dLLM and introduce the weighted direct discriminative optimization (WDDO) loss, again through importance sampling to represent $p_{*}({\bm{o}}|{\bm{q}})$,
$\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=-\underset{{\bm{\sigma}}}{\operatorname{\mathbb{E}}}\,\underset{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{\operatorname{\mathbb{E}}}\left[w({\bm{o}}|{\bm{q}};{\bm{\sigma}})\log\sigma\left(\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\right)+\log\sigma\left(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\right)\right],$
where $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ is the importance weight defined in ˜10.
@@ 4 Experiments
Model and baselines. We apply DMPO to LLaDA-8B-Instruct (Nie et al., 2025b), a state-of-the-art open-sourced, native masked dLLM that has not been post-trained with RL techniques. To clearly demonstrate the potential of DMPO, we follow an R1-Zero-like training recipe (Guo et al., 2025a; Liu et al., 2025c) and directly apply DMPO to the LLaDA model without first performing SFT on curated datasets. We refer to the model obtained via this pipeline as DMPO-LLaDA. We benchmark our method against a series of top-performing dLLM base models of comparable model size, such as Dream-Instruct (7B, Ye et al. (2025)), LLaDA-Instruct (8B, Nie et al. (2025b)), LLaDA-1.5 (8B, Zhu et al. (2025a)). Our main RL baseline is d1 (Zhao et al., 2025a), a state-of-the-art RL finetuning approach developed for dLLMs that combines both SFT and diffu-GRPO (an adapted version of GRPO). In the main result table Tab.˜1, DMPO-LLaDA uses the group weight baseline on GSM8K, MATH500, and Sudoku, while employing the individual weight baseline on Countdown.
@@
Experimental setups. We perform experiments on 4 different reasoning benchmarks: GSM8k (Cobbe et al., 2021), MATH500 (Lightman et al., 2023; Hendrycks et al., 2021), Sudoku (Arel, 2025), and Countdown (Pan et al., 2025). For all pretrained dLLM models, we evaluate the latest available checkpoints for each task. For d1, we reproduce their results following the exact guidelines provided. To ensure a fair comparison, we train DMPO-LLaDA on the same datasets as d1 for each task with rollouts generated using a fixed sequence length of $256$. All evaluations are conducted with zero-shot prompting using three different generation lengths: $128$, $256$, and $512$, following a similar practice as in Zhao et al. (2025a). See App.˜C for more details of experimental setups.
Figure 4: Rewards dynamics during training. DMPO consistently produces higher rewards than d1.
Figure 5: Effects of negative gradient insertion for DMPO.
@@
DMPO incentivizes superior reasoning capabilities. We report in Tab.˜1 the performance of DMPO together with that of the base model LLaDa-Instruct (8B), the models obtained by d1 post-training strategies, and other pretrained dLLM models. DMPO consistently outperforms both the LLaDA-Instruct baseline and the d1 models, achieving the best performance among the listed state-of-the-art dLLM models. Notably, DMPO achieves excellent gains over the LLaDA-Instruct baseline, with an accuracy improvement of an average of $\bm{+2.40\%}$ on GSM8K, $\bm{+3.0\%}$ on MATH500, $\bm{+55.8\%}$ on Countdown, $\bm{+16.96\%}$ on Sudoku. DMPO also demonstrates superior performance over d1, the current SOTA RL baseline for dLLM, especially on planning tasks, with an increase of $\bm{+42.9\%}$ on Countdown and $\bm{+11.8\%}$ on Sudoku. This underscores the overall effectiveness of DMPO for enhancing model reasoning capabilities.
DMPO consistently achieves higher rewards. In Fig.˜4, we present the reward dynamics of DMPO across training steps and compare with that of d1. DMPO consistently achieves higher reward values after an initial warm-up phase and ultimately discovers responses with higher reward values than d1, possibly due to its continuous exploration of the reward distribution landscape throughout the training process. While in the first $1,000$ steps, DMPO often produces lower reward values than d1, this is potentially due to the lack of the SFT phase before performing RL scaling. Moreover, we observe that the performance of DMPO does not appear to saturate after $4,000$ gradient steps, suggesting its greater potential than GRPO-type algorithms for dLLM fine-tuning.
@@
Weight baseline subtraction is crucial for small batch size training. We test the different choices presented for negative gradient insertion in Secs.˜3.3 and 3.4 when training on the Sudoku dataset with a small batch size, and the result is visualized in Fig.˜5. As is clear from the curves, when weight baseline subtraction is not employed, the model does not improve as training progresses. Across the three presented weight baseline choices for WDCE, the group weight baseline ˜13 consistently outperforms the other two choices. Weighted DDO promotes the fastest reward increase among the initial $1,200$ steps, but suffers from instability and collapse issues afterward.

### v2

@@ 3.3 Effective Training with Negative Gradient Insertion
We note that this issue does not arise when the batch size is sufficiently large for the following reason. When having a large batch of diverse responses that make up a good coverage of the sample space, despite having all positive weights, since the model cannot increase likelihood on all responses (as the probabilities sum up to $1$), the “bad” responses will be automatically and implicitly penalized due to not having larger weights than the “good” responses.
@@
When the batch size is small, the scenario is different as is illustrated in Fig. 3. In such a case, the model will tend to promote both “good” and “bad” responses due to the positive weights, and potentially penalize the likelihood of other unseen responses to maintain a valid probability distribution. This could be detrimental to achieving distribution matching, as these unseen responses may have high reward values and correspond to an undiscovered distribution mode.
@@
To address this issue, we inject negative gradient (Ren and Sutherland, 2025; Deng et al., 2025) by designing a weight baseline and subtract it from the obtained weights to facilitate an effective reinforcement on the good samples, i.e.,
$w_{(}{\bm{o}}|{\bm{q}};{\bm{\sigma}})=w({\bm{o}}|{\bm{q}};{\bm{\sigma}})-w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}}).$ (12)
@@
Group weight baseline
When the dLLM policy is close to optimal, the original log weight $\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ should behave approximately like constants for a group of different responses $\{{\bm{o}}^{(n)}\}_{1\leq n\leq N}$, leading to nearly uniform weight value for $\{w({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})\}_{1\leq n\leq N}$ after group softmax. We can thus choose the baseline as 11 to encourage convergence to this optimal situation:
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=1.$ (13)
Individual weight baseline
We can also consider the individual weight value of each response. For samples with smaller log weights, a stronger penalization is more desirable. A natural, adaptive way to design penalization strength is to use softmax over the log weights with negative reward: define $\ell_{-}({\bm{o}}|{\bm{q}};{\bm{\sigma}}):=-\frac{r({\bm{q}},{\bm{o}})}{\alpha}+\log\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$, and
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=\frac{N\exp(\ell_{-}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)}))}{\sum_{k}\exp(\ell_{-}({\bm{o}}^{(k)}|{\bm{q}};{\bm{\sigma}}^{(k)}))}.$ (14)
Note that this $w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ now corresponds to a bad target distribution given by $p_{*-}({\bm{o}}|{\bm{q}})\propto_{\bm{o}}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{-r({\bm{q}},{\bm{o}})/\alpha}$, which is tilted by the negative reward. The minus sign in the loss before $w_{\mathrm{base}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ means we want to drive the dLLM policy away from this bad distribution.
Model weight baseline
Finally, we can determine whether to promote or penalize specific responses by comparing $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ with the importance weight under the current model policy $p_{\theta}({\bm{o}}|{\bm{q}})$, which pushes the model further towards the optimal $p_{*}({\bm{o}}|{\bm{q}})$. Note that this does not incur additional computation overhead as we can estimate $\log p_{{\bar{\theta}}}({\bm{o}}|{\bm{q}})$ using negative ELBO 2, which is already computed in the WDCE loss. Define $\ell_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}}):=\log\frac{p_{{\bar{\theta}}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$, and let
$w_{\mathrm{base}}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)})=\frac{N\exp(\ell_{\theta}({\bm{o}}^{(n)}|{\bm{q}};{\bm{\sigma}}^{(n)}))}{\sum_{k}\exp(\ell_{\theta}({\bm{o}}^{(k)}|{\bm{q}};{\bm{\sigma}}^{(k)}))}.$ (15)
We remark that the group weight and model weight baselines 13 and 15 can also be interpreted as an approximate variance reduction. See Sec. B.3 for discussion.
@@
d1-LLaDA $75.28$ $81.40$ $84.38$ $30.00$ $36.60$ $40.80$ $34.38$ $26.56$ $30.47$ $21.97$ $11.04$ $8.69$
cGRPO-LLaDA $67.40$ $81.73$ $84.23$ $21.40$ $32.80$ $38.40$ $30.08$ $42.58$ $37.11$ $24.17$ $24.17$ $21.97$
DMPO-LLaDA (Ours) $74.83$ $82.41$ $\mathbf{85.22}$ $30.00$ $38.20$ $\mathbf{42.80}$ $\mathbf{67.19}$ $\mathbf{80.86}$ $82.81$ $\mathbf{32.76}$ $24.56$ $19.97$
DMPO-LLaDA-SFT (Ours) $\mathbf{80.06}$ $\mathbf{84.00}$ $84.09$ $\mathbf{31.80}$ $\mathbf{40.00}$ $41.20$ $54.69$ $67.19$ $77.34$ $25.20$ $\mathbf{25.73}$ $\mathbf{23.78}$
DMPO-LLaDA-1.5 (Ours) $77.56$ $82.71$ $84.61$ $30.20$ $36.60$ $41.00$ $59.77$ $79.30$ $\mathbf{83.20}$ $25.34$ $24.51$ $23.34$
@@ 3.4 Weighted Direct Discriminative Optimization
To explore the full potential of the distribution matching framework in 6, we also investigate other choices for the potential $\mathcal{F}$ beyond the cross-entropy. One particularly interesting objective is the following direct discriminative optimization (DDO) loss:
$\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=-\operatorname{\mathbb{E}}_{p_{*}({\bm{o}}|{\bm{q}})}\log\sigma\Big(\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\Big)-\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}})}\log\sigma\Big(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\Big),$ (16)
where $\sigma(t)=1/(1+\mathrm{e}^{-t})$. The global minimizer of 16 is also $p_{*}(\cdot|{\bm{q}})$, thus being a valid functional for distribution matching. For a more detailed justification, see Sec. B.4.
This is inspired by Zheng et al. (2025e), which proposed a GAN-like (Goodfellow et al., 2014) loss for SFT of vision models. One interesting trait of this objective is its natural incorporation of negative gradients for bad samples due to the GAN nature, as is shown in the analysis therein:
$\nabla_{\theta}\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=\sum_{\bm{o}}\sigma\Big(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\Big)\cdot\boxed{(p_{\theta}({\bm{o}}|{\bm{q}})-p_{*}({\bm{o}}|{\bm{q}}))}\nabla_{\theta}\log p_{\theta}({\bm{o}}|{\bm{q}}).$
From the expression, as the first term is always non-negative, and the boxed term applies a penalty for bad response $\bm{o}$}, thus providing a gradient direction for increasing $p_{\theta}({\bm{o}}|{\bm{q}})$. Leveraging this property, we adapt it for RL finetuning of dLLM and introduce weighted direct discriminative optimization (WDDO) loss, again dealing with $p_{*}$ through importance sampling:
$\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}}))=-\operatorname{\mathbb{E}}_{{\bm{\sigma}}}\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\Big[w({\bm{o}}|{\bm{q}};{\bm{\sigma}})\log\sigma\Big(\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\Big)+\log\sigma\Big(-\log\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}\Big)\Big],$
where $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ is the importance weight defined in 10.
Figure 4: Effects of negative gradient insertion on Sudoku.
Figure 5: Comparison of training dynamics on Countdown. $F$ is the frequency of sampling the buffer.
@@ 4 Experiments
We apply DMPO to LLaDA-8B-Instruct (Nie et al., 2025b), a SOTA dLLM not post-trained with RL techniques. To clearly demonstrate the potential of DMPO, we follow an R1-Zero-like training recipe (Guo et al., 2025a; Liu et al., 2025c) and directly apply DMPO to the LLaDA model without first performing SFT on curated datasets. We refer to the model obtained from this pipeline as DMPO-LLaDA. We benchmark our method against a series of dLLM base models of comparable size such as Dream-Instruct (7B, Ye et al. (2025)), LLaDA-Instruct (8B, Nie et al. (2025b)), and LLaDA-1.5 (8B, Zhu et al. (2025a)). Our main RL baseline is d1 (Zhao et al., 2025a), a SOTA dLLM RL finetuning approach that combines both SFT and diffu-GRPO (an adapted version of GRPO). We also compare with cGRPO, which was used to finetune a coding dLLM in Gong et al. (2025).
Experimental setups
@@
We explore four reasoning benchmarks: GSM8k, MATH500, Sudoku, and Countdown. For all pretrained dLLMs, we evaluate the latest available checkpoint. For d1 and cGRPO, we reproduce their results exactly following the provided guidelines. To ensure a fair comparison, we train DMPO-LLaDA on the same datasets as d1 for each task with rollouts generated using a fixed sequence length of $256$. All evaluations are conducted with zero-shot prompting using three generation lengths: $128$, $256$, and $512$, following a similar practice as in Zhao et al. (2025a).
DMPO incentivizes superior reasoning capabilities
@@
We report in Tab. 1 the performance of DMPO together with that of the base model LLaDA-Instruct, LLaDA-1.5, the models obtained by d1 and cGRPO post-training strategies, and other pretrained dLLMs. DMPO consistently outperforms both the LLaDA-Instruct baseline and the d1/cGRPO models, achieving the best performance among the listed SOTA dLLMs. Notably, it achieves excellent gains over LLaDA-Instruct, with an accuracy improvement of an average of $\bm{+2.40\%}$ on GSM8K, $\bm{+3.00\%}$ on MATH500, $\bm{+59.38\%}$ on Countdown, $\bm{+16.96\%}$ on Sudoku. DMPO also demonstrates superior performance over d1, the current SOTA RL baseline for dLLMs, especially on planning tasks, with an increase of $\bm{+46.48\%}$ on Countdown and $\bm{+11.86\%}$ on Sudoku. This underscores the overall effectiveness of DMPO for enhancing model reasoning capabilities.
Weight baseline subtraction is crucial for small batch size training
@@
We test the different choices for negative gradient insertion in Secs. 3.3 and 3.4 when training on the Sudoku dataset with a small batch size, and the result is visualized in Fig. 4. As shown by the curves, without weight baseline subtraction, the model does not improve as training progresses. All the proposed weight baselines in 13, 14 and 15 effectively increase the reward value during training. WDDO achieves the fastest reward increase during the initial $1\mathrm{k}$ steps but suffers from instability afterwards.
DMPO benefits from other means of post-training techniques
To showcase the robustness and general efficacy of DMPO, we apply it to LLaDA-SFT and LLaDA-1.5. LLaDA-SFT is obtained by performing SFT of LLaDA-Instruct (8B) on s1k (Muennighoff et al., 2025), a dataset of $1\mathrm{k}$ examples of high-quality reasoning questions with distilled reasoning traces from Gemini Thinkin;. LLaDA-1.5 is obtained by performing DPO on 350K preference pairs covering a wide range of topics such as writing and reasoning. We then apply DMPO to these base models to obtain DMPO-LLaDA-SFT and DMPO-LLaDA-1.5, with performance reported in Tab. 1. DMPO continues to deliver performance gains for post-trained models, with consistent and significant accuracy improvements over base models, especially at generation lengths of $128$ and $256$ for the math reasoning datasets, with $\bm{+4.78\%}$ and $\bm{+2.60\%}$ on GSM8K, $\bm{+1.80\%}$ and $\bm{+3.40\%}$ on MATH500 compared with d1. This underscores that DMPO is a powerful method that integrates smoothly with existing solutions.
DMPO enables efficient and fast training
Due to its off-policy and forward nature, DMPO achieves considerable training acceleration compared with GRPO-type methods. In Fig. 5, we compare head-to-head the training dynamics of diffu-GRPO, DMPO with ROAR sampler, and DMPO with Fast-dLLM (an approximate KV-cache mechanism enabled, confidence-based heuristic sampler for dLLMs, from Wu et al. (2025)) on Countdown under the same amount of training compute. Due to its off-policy nature, DMPO enables heavy reuse of each sampled buffer of rollouts and achieves a sample efficiency $\bm{2\sim 3\times}$ that of diffu-GRPO. Regarding training-compute efficiency, as a forward-loss-based algorithm, DMPO enjoys a flexible choice of rollout sampler. With fast-dLLM, DMPO gains an acceleration of up to $\bm{8\times}$ per rollout sampling, and achieves the same level of reward as d1 with only $\bm{31\%}$ of the training budget ($1.8$ PF-days 22211 PF-day $8.64\times 10^{19}$ floating point operations. v.s. $5.78$ PF-days). This empirical evidence emphasizes that DMPO is not only an effective algorithm but also highly sample- and compute-efficient.
DMPO exhibits stable training despite highly stale data
As is evident from Figs. 6 and 5, DMPO enjoys a largely stable dynamics despite using up to $24\times$ stale data (which means $24$ parameter updates on the same batch of rollouts), without suffering from high variance of importance sampling. While this seems to contradict the general belief that on-policy learning beats off-policy learning for LLM RL, we argue that this is not the case, because the off-policy in DMPO is inherently different from that used in diffu-GRPO or GRPO. Note that the latter considers the importance weight of the form $\frac{\pi_{\theta}}{\pi_{\text{old}}}$, which inevitably diverges as the number of parameter updates on $\theta$ increases. However, DMPO uses importance weight of the form $\frac{p_{*}}{p_{\mathrm{old}}}$, which is independent of the current policy model $\pi_{\theta}$ and remains stable over a long horizon of training, enabling the use of a low buffer sampling frequency and highly stale rollouts without sacrificing performances. Moreover, DMPO adopts sequence-level importance sampling, in contrast to the token-level importance sampling used in diffu-GRPO or cGRPO, thereby providing an additional layer of stability. This advantage is also discussed in depth in Group Sequence Policy Optimization (GSPO, Zheng et al. (2025a)), which similarly considers sequence-level importance sampling.
Additional results and discussion can be found in Sec. C.3.